# Paying and Persuading

### v4

@@ 4.1 Feasible Optimality
Rearranging and taking expectations, an action $a$ is feasibly optimal if for all $\mu$
$\mathbb{E}_{\mu}\left[\frac{v(a^{\prime},\theta)-v(a,\theta)}{u(a^{\prime},\theta)-u(a,\theta)}\right]>-k$
whenever it is well defined. Feasibly optimal actions are named exactly because they are the only actions which can appear on the path of play in any optimal contract.
@@
Let $a_{1}=\operatorname*{\arg\!\max}_{a\in A}v(a)$ be Sender’s favorite action and suppose $v(a_{1})-v(a)>k$ for all $a\in A\setminus\{a_{1}\}$. Then at every $h^{\xi}$, $\alpha(h^{\xi})(\mu)=\delta_{a_{1}}$.
Corollary 1 thus gives a unique prediction when Sender has state-independent preferences and $k$ is sufficiently low. While straightforward given Proposition 3, it will simplify the characterization of the optimal contract in the leading example (Proposition 10). In general, the feasibly optimal set will also play a role in characterizing the set of beliefs which can arise on the path of play at any optimum as well, as I next characterize.
@@ 4.2 Effectiveness Ratios
Let $\mathcal{E}(k)=\{\mu\in\Delta(\Theta)^{o}:E(\mu)>-k\}$ be the set of all beliefs whose effectiveness ratio is strictly greater than k-k. Here, $a^{*}(\mu,\mathbf{0})\in\operatorname*{\arg\!\max}_{a\in A}[u(a,\theta)]$, abusing notation to refer to the correspondence as a function when there is no loss of ambiguity.
@@
Suppose full-information is strictly optimal for Receiver. Then for any optimal $\sigma$ and $(h^{\xi},\mu,a)\in\mathbb{P}^{\sigma}$, if $t(h^{\xi};\sigma)(\mu,a)>0$, then $\text{supp}(\rho(h^{\zeta};\sigma))\subset\left(\mathcal{E}(k)^{c}\right)$ for all $h^{\zeta}\succ(h^{\xi},\mu,a)$.
@@
The proof follows from some straightforward computations (these can be found in the PROOF OF PROPOSITION 5 ) I interpret Proposition 5 as a comparative statics result: as transfers become relatively more expensive, the set of beliefs that are ruled out strictly expands since $\text{supp}(\rho(h^{\zeta};\sigma))\subset\mathcal{E}(k)^{c}$, so that only those beliefs which are “more extremal” remain consistent with optimality after payments. As $k\to\infty$, the degenerate beliefs (i.e. the extreme points of the probability simplex) become the only beliefs consistent with payments. This is because Sender can always, in lieu of payments, reveal the state and have Receiver take their favorite action, instead of using transfers to change Receiver actions.
@@ 4.3 Incentivizability
One way to interpret Theorem 1 is that there exists some history-dependent stopping time $T_{\sigma}^{*}$ such that at each history, before $T_{\sigma}^{*}$, only (dynamic) informational incentives are used to motivate Receiver, while after $T_{\sigma}^{*}$, only transfers are used as a way to motivate Receiver. Theorem 1 leaves open the question of whether Sender finds it strictly profitable to sequence information and transfers. In particular, one might worry that the time structure of optimal information and transfers holds only because all combinations of information and transfers are optimal, and hence Theorem 1 only isolates one such contract. However, this is not the case—this subsection identifies conditions that guarantee dynamic informational incentives are used strictly before transfers in any optimum. Moreover, these conditions are natural: they are satisfied whenever repeated static persuasion is not optimal in the class of all stationary strategies. [^14] In Appendix C: Incentivizability with Continuous Actions , I show that in sufficiently continuous problems, the regularity conditions identified here hold for a broad class of economic primitives. Proposition 6 thus highlights the importance of (strictly) frontloading informational incentives before paying Receiver to motivate them.
I suppose throughout this subsection the state is i.i.d. and normalize payoffs so that $\underline{U}(\mu_{0})=0$. I adopt the following definition for when, across all optimum, informational incentives must take strict precedence over monetary ones.
@@
What are the conditions under which transfers are strictly backloaded at every optimum? Two conditions are clearly necessary. First, it must be that the Pareto frontier is not flat, i.e. $V_{+}^{\prime}(0,\mu_{0})\neq-k$. If this were the case, then Lemma 4 would imply transfers and information were exactly perfect substitutes, and hence any combination of (pulled back) information and transfers would be part of an optimal way to move along the Pareto frontier. Second, there must be some role for dynamic incentives, so that a stationary optimum where Receiver always takes $a^{*}(\mu,\mathbf{0})$ is not always optimal. If this were the case, then there would be no role for dynamic incentives, either through transfers or information. The next result shows these two necessary conditions exactly characterize when transfers are strictly backloaded.
@@
The proof can be found in the PROOF OF PROPOSITION 6 Nontriviality is the easier of the two conditions to satisfy, and in particular will hold so long as $k$ is sufficiently large (i.e. transfers are expensive enough that information is “cheaper” at a first pass). Incentivizability is the less obvious of the two conditions to be satisfied, since it requires control over every solution to the dynamic contracting problem, which is itself an endogenous equilibrium object. I conclude this section by giving an easier to check sufficient condition for incentivizability.
For the statement of the next result, let $\mathcal{V}^{*}(\mu_{0},0)$ be the value to the static problem, and $\text{cav}(V_{0})(\mu_{0})$ to be the solution to the static problem without transfers. Note that $\mathcal{V}^{*}(\mu_{0},\delta)>\text{cav}(V_{0})(\mu_{0})$ if and only if transfers are useful to Sender in the static one-shot problem. Proposition 7 shows that so long as transfers are useful in the static one-shot problem, then the problem is incentivizable.
@@
Note Proposition 7 gives a condition for incentivizability for any $\delta>0$, as a function of behavior at $\delta=0$. A formal proof is in PROOF OF PROPOSITION 7 To see why the result is true, note that if the problem is not incentivizable, then there is a stationary optimum where $a=a^{*}(\mu,\mathbf{0})$ on the path of play. But this solution, which is feasible in the information design problem without transfers, must be optimal for the dynamic contract where $\delta=0$ as well, since utility promises and intertemporal transfers are never used. Hence $\mathcal{V}^{*}(\mu_{0},0)=\text{cav}(V_{0})(\mu_{0})$.
One objection to Proposition 7 is that verifying whether transfers are used in the static problem might itself be a computationally difficult exercise, and hence Proposition 7 simply punts one relatively inscrutable condition (incentivizability) into a different one. Luckily, this is not the case; in the next section, I develop a simple algorithm that verifies whether transfers are used in the joint static problem when $\delta=0$ by looking at “maximally indifferent” beliefs.
@@ 5.1 The Myopic Case
Consider the case where $\delta=0$ so Sender and Receiver are fully myopic. In this case, it is without loss of generality to focus on only the first period, and so Sender’s strategy is simply a transfer function $t:\Delta(\Theta)\times A\to\mathbb{R}_{+}$ and an experiment $\rho\in\Delta(\Delta(\Theta))$ which is $\mu_{0}$-Bayes plausible. For any transfer rule $t$, define Receiver’s best response (breaking ties in favor of Sender) as
@@
where we appeal to Lemma 1 to simplify the space of transfers. As before, let $a^{*}(\mu,\mathbf{0})$ to be Receiver’s optimal action in the absence of transfers.
@@
$\max_{\rho\in\Delta_{\mu_{0}}(\Delta(\Theta)),t}\left\{\mathbb{E}_{\mu}[v(a^{*}(\mu,t),\theta)-kt(\mu,a))]\right\}.$
@@
$V^{*}(\mu_{0})=\text{cav}|_{\mu_{0}}\left(\max_{t}\mathbb{E}_{\mu}[v(a^{*}(\mu,t),\theta)-kt(\mu,a))]\right)$
@@
Proposition 8 is proven in the Appendix F: Proof of Proposition 8 Note even though the transfer-augmented function has Receiver take surplus maximizing actions at every belief, Sender need not fully reveal the state, in contrast to standard intuition. This is because of the limited liability constraint: there may exist beliefs where Sender would prefer to charge Receiver if they could at the surplus maximizing action, but cannot. Consequently, Sender may do better by pooling that belief with one where Receiver is given less surplus (i.e. made maximally indifferent) to persuade Receiver to take actions which are more beneficial to Sender. Hence the ability to persuade can be strictly useful even when full information is ex-post efficient so long as Sender faces a limited liability constraint.
@@
For any finite $A$, $\text{cav}(V_{0}(\mu))$ is piecewise affine.
@@ 5.2 The Patient Ergodic Case
$\Gamma(u,\mu)=\left\{(\gamma,m):\mathbb{E}_{\gamma}[u(a,\theta)]+m\geq u\text{ and }\gamma|_{\Theta}=\mu\right\}$
@@
is the value of Receiver’s outside option at the ergodic distribution $\mu^{\infty}$. $\Gamma(\underline{U}(\mu^{\infty}))$ plays an important role in bounding the value of persuasion in Proposition 9. To operationalize the bound, I need one more definition.
@@
Moreover if there is an optimal stationary $\sigma^{*}$, then the inequality holds with equality.
The formal proof can be found in the PROOF OF PROPOSITION 9 The intuition uses the exponential concentration of distributions to their steady state, regardless of the initial distribution to show that as $\delta\to 1$, “essentially” only Receiver incentive compatibility at the ergodic steady-state matters. Thus, if (and only if) Receiver prefers the induced joint distribution to their no-information payoff, then some stationary distribution leading to $\gamma$ is implementable. The upper bound can be seen as implying the individual rationality constraint characterizes the value of optimal stationary contracts [^17] With transfers, these are often those which trace out the efficient set when agents have relational incentives; see Levin (2003) and Kolotilin and Li (2021). exactly. However, outside of special cases, knowing when there is an optimal stationary contract can be difficult.
@@
Bounding Sender’s optimal payoff can also shed qualitative light on the nature of dynamic persuasion away from Receiver’s individual rationality constraint. First, it implies that a sufficient condition for Sender to strictly benefit from dynamics is that Receiver’s incentive compatibility constraint is not binding.
@@
Let $\gamma$ be the coupling concavifying $V^{t}(\mu_{0})$. If $\mathbb{E}_{\gamma}[u(a,\theta)]>0$, then $\mathcal{V}^{*}(\mu_{0},\delta)>\text{cav}(V^{t}(\mu_{0}))$ unless $\text{cav}(V^{t}(\mu_{0}))=\mathbb{E}_{\mu_{0}}[\max_{a\in A}[v(a,\theta)]]$.
Corollary 3 contrasts with Theorem 1 of Ely (2017) and highlights the value of the obedience-based characterization in Proposition 9 over the recursive concavification approach. In particular, unlike Ely (2017), I can give easy to verify conditions under which Sender benefits from the ability to dynamically persuade; moreover, these conditions hold for all $\delta$, and are independent of the underlying Markov chain, and hence are easy to verify.

### v5

@@ 4.1 Feasible Optimality
Rearranging and taking expectations, an action $a$ is feasibly optimal if, for all $\mu$, the marginal rate of substitution between Sender and Receiver preferences is at least k-k; that is, for all beliefs $\mu$,
$\mathbb{E}_{\mu}[v(a^{\prime},\theta)-v(a,\theta)]>-k{\mathbb{E}_{\mu}[u(a^{\prime},\theta)-u(a,\theta)]}$
Feasibly optimal actions are named exactly because they are the only actions which can appear on the path of play in any optimal contract.
@@
Let $a_{1}=\operatorname*{\arg\!\max}_{a\in A}v(a)$ be Sender’s favorite action and suppose $v(a_{1})-v(a)>k\max_{\theta}\{u(a,\theta)-u(a_{1},\theta)\}$ for all $a\in A\setminus\{a_{1}\}$. Then at every $h^{\xi}$, $\alpha(h^{\xi})(\mu)=\delta_{a_{1}}$.
Corollary 1 thus gives a unique prediction when Sender has state-independent preferences and $k$ is sufficiently low. While straightforward given Proposition 3, it will simplify the characterization of the optimal contract in the leading example (Theorem 2). In general, the feasibly optimal set will also play a role in characterizing the set of beliefs which can arise on the path of play at any optimum as well, as I next characterize.
@@ 4.2 Effectiveness Ratios
Let $\mathcal{E}(k)=\{\mu\in\Delta(\Theta)^{o}:E(\mu)>-k\}$ be the set of all beliefs whose effectiveness ratio is strictly greater than k-k. Here, $a^{*}(\mu,\mathbf{0})\in\operatorname*{\arg\!\max}_{a\in A}\mathbb{E}_{\mu}[u(a,\theta)]$, abusing notation to refer to the correspondence as a function when there is no loss of ambiguity.
@@
Suppose full-information is strictly optimal for Receiver. Then for any optimal $\sigma$ and $(h^{\xi},\mu,a)\in\mathbb{P}^{\sigma}$, if $t(h^{\xi};\sigma)(\mu,a)>0$, then $\text{supp}(\rho(h^{\zeta};\sigma))\subset\mathcal{E}(k)^{c}$ for all $h^{\zeta}\succ(h^{\xi},\mu,a)$.
@@
The proof follows from some straightforward computations (these can be found in PROOF OF PROPOSITION 5 ) I interpret Proposition 5 as a comparative statics result: as transfers become relatively more expensive, the set of beliefs that are ruled out strictly expands since $\text{supp}(\rho(h^{\zeta};\sigma))\subset\mathcal{E}(k)^{c}$, so that only those beliefs which are “more extremal” remain consistent with optimality after payments. As $k\to\infty$, the degenerate beliefs (i.e. the extreme points of the probability simplex) become the only beliefs consistent with payments. This is because Sender can always, in lieu of payments, reveal the state and have Receiver take their favorite action, instead of using transfers to change Receiver actions.
@@ 4.3 Incentivizability
One way to interpret Theorem 1 is that there exists some history-dependent stopping time $T_{\sigma}^{*}$ such that at each history, before $T_{\sigma}^{*}$, only (dynamic) informational incentives are used to motivate Receiver, while after $T_{\sigma}^{*}$, only transfers are used as a way to motivate Receiver. Theorem 1 leaves open the question of whether Sender finds it strictly profitable to sequence information and transfers. In particular, one might worry that the time structure of optimal information and transfers holds only because all combinations of information and transfers are optimal, and hence Theorem 1 merely isolates one such contract. However, this is not the case—this subsection identifies conditions that guarantee dynamic informational incentives are used strictly before transfers in any optimum. Moreover, these conditions are natural: they are satisfied whenever repeated static persuasion is not optimal in the class of all stationary contracts. [^16] In Appendix C: Incentivizability with Continuous Actions , I show that in sufficiently continuous problems, the regularity conditions identified here hold for a broad class of economic primitives. Proposition 6 thus highlights the importance of (strictly) frontloading informational incentives before paying Receiver to motivate them.
I suppose throughout this subsection the state is i.i.d. and normalize payoffs so that $\underline{U}(\mu_{0})=0$. [^17] This assumption is made for expositional convenience and is not necessary for Proposition 6. I adopt the following definition for when, across all optimum, informational incentives must take strict precedence over monetary ones.
@@
What are the conditions under which transfers are strictly backloaded at every optimum? Two conditions are clearly necessary. First, it must be that the Pareto frontier is not flat, i.e. $V_{+}^{\prime}(0,\mu_{0})\neq-k$. If this were the case, then Lemma 4 would imply transfers and information were exactly perfect substitutes, and hence any combination of (pulled back) information and transfers would be part of an optimal way to move along the Pareto frontier. Second, there must be some role for dynamic incentives, so that a stationary optimum where Receiver always takes $a^{*}(\mu,\mathbf{0})$ is not always optimal. If this were the case, then there would be no role for dynamic incentives, either through transfers or information. The next result shows these two clearly necessary conditions are in, fact, exactly sufficient: they characterize when transfers are strictly backloaded.
@@
Proposition 6 strengthens Theorem 1 in the following way: it provides a history-dependent stopping time $T_{\sigma}^{*}$, before which only informational incentives are used to motivate Receiver (and indeed they are used), and after which transfers can potentially be used. The proof can be found in the PROOF OF PROPOSITION 6 Nontriviality is the easier of the two conditions to satisfy, and in particular will hold so long as $k$ is sufficiently large (i.e. transfers are expensive enough that information is “cheaper” at a first pass). Incentivizability is the less obvious of the two conditions to be satisfied, since it requires control over every solution to the dynamic contracting problem, which is itself an endogenous equilibrium object. I conclude this section by giving an easier to check sufficient condition for incentivizability.
For the statement of the next result, let $\mathcal{V}^{*}(\mu_{0},0)$ be the value to the static problem, and $\text{cav}(V_{0})(\mu_{0})$ to be the solution to the static problem without transfers. Note that $\mathcal{V}^{*}(\mu_{0},0)>\text{cav}(V_{0})(\mu_{0})$ if and only if transfers are useful to Sender in the static one-shot problem. Proposition 7 shows that so long as transfers are useful in the static one-shot problem, then the problem is incentivizable.
@@
Note Proposition 7 gives a condition for incentivizability for any $\delta>0$, as a function of behavior at $\delta=0$. A formal proof is in PROOF OF PROPOSITION 7 To see why the result is true, note that if the problem is not incentivizable, then there is an optimum where $a=a^{*}(\mu,\mathbf{0})$ always on the path of play. But this solution, which is feasible in the information design problem without transfers, must be optimal for the dynamic contract where $\delta=0$ as well, since utility promises and intertemporal transfers are never used. Hence $\mathcal{V}^{*}(\mu_{0},0)=\text{cav}(V_{0})(\mu_{0})$.
One objection to Proposition 7 is that verifying whether transfers are used in the static problem might itself be a computationally difficult exercise, and hence Proposition 7 simply punts one relatively inscrutable condition (incentivizability) into a different one. Luckily, this is not the case; in Section 5.1, I develop a simple algorithm that verifies whether transfers are used in the joint static problem when $\delta=0$ by looking at “maximally indifferent” beliefs.
@@
4.4 Expensive Transfers
The previous subsection gave conditions under which periods of pure information design are strictly preferred to using transfers in some periods of every optimal contract; that is, there are periods where dynamic information is strictly preferred to transfers as a motivating tool. This shows that optimal contracts consists of two nontrivial phases—one with only information, and one including transfers. However, it does not show every optimal contract must feature transfers as a last resort. In this subsection, I show that when transfers become sufficiently expensive (in the sense that $k$ is sufficiently large), then every optimal contract must feature transfers as a last resort. Formally,
Definition 7.
Transfers are always a last resort if for every optimal contract $\sigma^{*}$, transfers are a last resort at $\sigma^{*}$.
Proposition 8.
Fix any $\{u,v,\delta\}$. There exists some $\bar{k}\in\mathbb{R}_{+}$ such that for all $k>\bar{k}$, transfers are always a last resort at primitives $\{u,v,\delta,k\}$.
A formal proof is given in PROOF OF PROPOSITION 8 The intuition behind the proof is best read in contrast to the proof of Theorem 1. Recall the pullback lemma (Lemma 4) states that it is always possible to jointly provide information and transfers in the future at a rate of at most k-k. However, in the absence of transfers, there is a rate of providing future incentives using only information that depends only on $(u,v,\delta)$. So long as $k$ is larger than some bound on this rate, then the pullback lemma never needs to operate, since providing information as an incentive is always cheaper than jointly paying and persuading. Thus, every optimal contract will feature transfers as a last resort, using an argument akin to the squeezing lemma.
Three remarks are in order. First, the result is more subtle than simply stating that as $k\to\infty$, transfers are never used. For example, in the setting of Corollary 1 with high stakes (i.e. $v(a_{1})-v(a)$ is large), transfers will eventually be used on the path of play to ensure a constant action is induced even as they get more expensive. Hence transfers will be a last resort at every optimum as $k$ gets large, but will still be used on-path. Second, these conditions are complementary to, but do not subsume, the analysis in the previous two sections. In particular, the previous analysis held for any choice of primitives $\{u,v,\delta,k\}$, while this new analysis holds in the limit as $k\to\infty$. Finally, the result contrasts naturally with Theorem 1; while I require here that $k$ is large enough that the pullback argument is essentially trivial, Theorem 1 holds regardless of how cheap transfer are. This perhaps highlights the surprising nature of Theorem 1, as transfers are used as a secondary incentive instrument to information even as they become arbitrarily cheap.
@@ 5.1 The Myopic Case
Consider the case where $\delta=0$ so Sender and Receiver are fully myopic. In this case, it is without loss of generality to focus on only the first period, and so Sender’s contract is simply a transfer function $t:\Delta(\Theta)\times A\to\mathbb{R}_{+}$ and an experiment $\rho\in\Delta(\Delta(\Theta))$ which is $\mu_{0}$-Bayes plausible. For any transfer rule $t$, define Receiver’s best response (breaking ties in favor of Sender) as
@@
where I appeal to Lemma 1 to simplify the space of transfers. As before, let $a^{*}(\mu,\mathbf{0})$ to be Receiver’s optimal action in the absence of transfers.
@@
$\max_{\rho\in\Delta_{\mu_{0}}(\Delta(\Theta)),t}\left\{\mathbb{E}_{\rho}\left[\mathbb{E}_{\mu}[v(a^{*}(\mu,t),\theta)-kt(\mu,a^{*}(\mu,t)))]\right]\right\}.$
@@
$V^{*}(\mu_{0})=\text{cav}|_{\mu_{0}}\left(\max_{t}\mathbb{E}_{\mu}[v(a^{*}(\mu,t),\theta)-kt(\mu,a^{*}(\mu,t)))]\right)$
@@
Proposition 9 is proven in PROOF OF PROPOSITION 9 Note even though the transfer-augmented function has Receiver take surplus maximizing actions at every belief, Sender need not fully reveal the state, in contrast to standard intuition. This is because of the limited liability constraint: there may exist beliefs where Sender would prefer to charge Receiver if they could at the surplus maximizing action, but cannot. Consequently, Sender may do better by pooling that belief with one where Receiver is given less surplus (i.e. made maximally indifferent) to persuade Receiver to take actions which are more beneficial to Sender. Hence the ability to persuade can be strictly useful even when full information is ex-post efficient so long as Sender faces a limited liability constraint.
@@
For any finite $\Theta,A$, $\text{cav}(V_{0}(\mu))$ is piecewise affine.
@@ 5.2 The Patient Ergodic Case
$\Gamma(u,\mu)=\left\{(\gamma,m)\in\Delta(A\times\Theta)\times\mathbb{R}_{+}:\mathbb{E}_{\gamma}[u(a,\theta)]+m\geq u\text{ and }\gamma|_{\Theta}=\mu\right\}$
@@
is the value of Receiver’s outside option at the ergodic distribution $\mu^{\infty}$. $\Gamma(\underline{U}(\mu^{\infty}),\mu^{\infty})$ plays an important role in bounding the value of persuasion in Proposition 10. To operationalize the bound, I need one more definition.
@@
Moreover if there exists $\bar{\delta}$ such that there is an optimal stationary $\sigma^{*}$ for all $\delta>\bar{\delta}$, then the inequality holds with equality.
The formal proof can be found in the PROOF OF PROPOSITION 10 The intuition uses the exponential concentration of distributions to their steady state, regardless of the initial distribution to show that as $\delta\to 1$, “essentially” only Receiver individual rationality at the ergodic steady-state matters. Thus, if (and only if) Receiver prefers the induced joint distribution to their no-information payoff, then some stationary distribution leading to $\gamma$ is implementable. The upper bound can be seen as implying the individual rationality constraint characterizes the value of optimal stationary contracts [^20] With transfers, these are often those which trace out the efficient set when agents have relational incentives; see Levin (2003) and Kolotilin and Li (2021). exactly. However, outside of special cases, knowing when there is an optimal stationary contract can be difficult.
@@
Bounding Sender’s optimal payoff can also shed qualitative light on the nature of dynamic persuasion away from Receiver’s individual rationality constraint. First, it implies that a sufficient condition for Sender to strictly benefit from dynamics is that Receiver’s individual rationality constraint is not binding.
@@
Let $\gamma$ be the coupling concavifying $V^{t}(\mu_{0})$. If $\mathbb{E}_{\gamma}[u(a,\theta)]>\underline{U}(\mu^{\infty})$, then $\lim\limits_{\delta\to 1}\mathcal{V}^{*}(\mu_{0},\delta)>\text{cav}(V^{t}(\mu_{0}))$ unless $\text{cav}(V^{t}(\mu_{0}))=\mathbb{E}_{\mu_{0}}[\max_{a\in A}[v(a,\theta)]]$.
Corollary 3 contrasts with Theorem 1 of Ely (2017) and highlights the value of the obedience-based characterization in Proposition 10 over the recursive concavification approach. In particular, unlike Ely (2017), I can give easy to verify conditions under which Sender benefits from the ability to dynamically persuade; moreover these conditions are independent of the underlying Markov chain, and hence are easy to verify.